OpenStudy (anonymous):

Solve the system of equations algebraically. 4x-2y=8 2x-4y=4

OpenStudy (compassionate):

There are two methods. Substitution and Elimination. Which one do you think is the best to use?

OpenStudy (anonymous):

I was trying elimination but need help

OpenStudy (compassionate):

Okay, what step are you having trouble with?

OpenStudy (anonymous):

Okay so I know I have to multiply equations by values to make the coefficients of that variable opposites. I'm not getting the right answer though, so I don't know if I'm multiplying it by the wrong number

OpenStudy (compassionate):

4x-2y=8 2x-4y=4 Okay, so you want to cancel the terms. We can cancel either x or y. But, looking at this, lets cancel the y term first -2(4x-2y=8) -- If we multiply this equation by -2, y will cancel. 2x-4y=4 -8x + 4x = -16 2x - 4y = 4 Add the equations -6x = -12 Divide x = 2

OpenStudy (anonymous):

I understand now. thank you!